# Classical origins of Landau-incompatible transitions

### v1

@@ Abstract
Continuous phase transitions where symmetry is spontaneously broken are ubiquitous in physics and often found between ‘Landau-compatible’ phases where residual symmetries of one phase are a subset of the other. However, continuous ‘deconfined quantum critical’ transitions between Landau-incompatible symmetry-breaking phases are known to exist in certain quantum systems, often with anomalous microscopic symmetries. In this paper, we investigate the need for such special conditions. We show that Landau-incompatible transitions can be found in a family of well-known classical statistical mechanical models with anomaly-free on-site microscopic symmetries, introduced by José, Kadanoff, Kirkpatick and Nelson (Phys. Rev. B 16, 1217). The models are labeled by a positive integer $Q$ and constructed by a deformation of the 2d classical XY model, defined on any lattice, with an on-site potential that preserves a discrete $Q$-fold spin rotation and reflection symmetry. For a range of temperatures, even $Q$ models exhibit two Landau-incompatible partial symmetry-breaking phases and a direct transition between them for $Q\geq 4$. Characteristic features of Landau-incompatible transitions are easily seen, such as enhanced symmetries and melting of charged defects. For odd $Q$, and corresponding temperature ranges, two regions of a single partial symmetry-breaking phase are obtained, split by a stable ‘unnecessary critical’ line. We present quantum models with anomaly-free symmetries that also exhibit similar phase diagrams.
@@
Spontaneous symmetry breaking (SSB) underpins several important physical phenomena, from the development of long-range orders in matter to endowing mass to fundamental particles [1]. The simplest setting for SSB is when a phase of matter, classical or quantum, with a vacuum invariant under a symmetry group $G$ undergoes a phase transition to produce multiple vacua, each of which preserves only a subset of the original symmetries $H\subset G$. If such a phase transition is continuous, it can be described within the Landau-Ginzburg-Wilson-Fisher (LGWF) framework using a local order parameter field. About twenty years ago, the nature of exotic direct transitions between incompatible SSB quantum phases, where the vacuum symmetries of neither phase could be identified as a subset of the other, was investigated in two-dimensional quantum systems [2]. Although such transitions had appeared in earlier studies [3], Ref. [2] recognized the importance of the fact that they could not be framed within the LGWF paradigm in terms of order parameter fields. Instead, they were naturally formulated using gauge fields, which are hidden from sight in the ordered phases but appear at the transition. This prompted the moniker ‘deconfined quantum criticality’ (DQC) [2, 4].
@@
Figure 1: Phase diagrams for the JKKN Hamiltonian shown in Eq. 1 for $Q\geq 2$ and fixed $\gamma_{2}<0$. The arrows represent the different vacua characterized by the expectation value $\langle\theta_{j}\rangle$. The residual symmetry for each vacuum is shown following the notation in Eq. 9. For even $Q\geq 4$, the red line along $\gamma_{1}=0$, represents a direct transition between Landau-incompatible partial symmetry-breaking (p-SSB) phases. For odd $Q\geq 3$, it represents an unnecessary critical transition. The labels Ising, Ashkin-Teller (AT), 3-state Potts, and compactified free boson ($c=1$) indicate the conformal field theory describing the transition.
@@
What physical settings can give rise to such Landau-incompatible transitions? Descriptions using non-linear sigma models [5] and low-dimensional examples [6, 7] have made it clear that deconfined gauge fields are not necessary. However, in most known examples, DQC appears in quantum systems when microscopic symmetries are anomalous [8, 9]. In this paper, we show that most of these conditions are not necessary and Landau-incompatible transitions can, in fact, be found even in ordinary classical statistical mechanical systems with anomaly-free symmetries. We demonstrate this using a well known family of models introduced by José, Kadanoff, Kirkpatrick and Nelson (JKKN) [10] obtained by perturbing the 2d classical XY model by on-site potentials labeled by a positive integer $Q$. For even $Q\geq 4$, the phase diagram consists of a direct Landau-incompatible phase transition between two partial symmetry-breaking phases. This transition displays all notable characteristics of DQC, including the appearance of an enhanced symmetry that rotates between the order parameters of the Landau-incompatible phases and the melting of charged defects. Interestingly, neutral defects and charges also independently exist and can condense to produce Landau-compatible phase transitions. For odd $Q$, we find the appearance of an exotic second-order ‘unnecessary critical’ line separating two regions of single phase that does not represent a genuine phase transition [11, 12] but is nevertheless stable. To our knowledge, this is the first case of unnecessary criticality identified in a classical model. Finally, we also present a family of quantum models with anomaly-free symmetries that exhibit similar phase diagrams. Altogether, our results show that exotic transitions beyond the LGWF paradigm are more abundant than previously believed.
Models and phase diagram:
@@
Let us consider the JKKN model [10] which is a classical statistical mechanical system of planar rotors $\theta_{j}\sim\theta_{j}+2\pi$, located on the vertices of any two-dimensional lattice with Hamiltonian
$H=-\sum_{\langle j,k\rangle}\cos\left(\theta_{j}-\theta_{k}\right)-\sum_{\ell=1,2,\ldots}\gamma_{\ell}\sum_{j}\cos\left(\ell Q\theta_{j}\right);$ (1)
@@
here $j$ labels the vertices and $\langle j,k\rangle$ the edges of the lattice. The relevant symmetry group of the system is generated by a rotation ${\mathcal{R}}$ and a reflection ${\mathcal{T}}$ acting as
@@
and thus form the non-abelian dihedral group $D_{2Q}\cong{\mathbb{Z}}_{Q}\rtimes\mathbb{Z}_{2}$. To determine the phase diagram of Eq. 1, we will study the partition function
[…]
varying $\gamma_{1}$ and the temperature $T=\beta^{-1}$ while keeping all other terms fixed, the most important being $\gamma_{2}$. We will be interested in the phase diagrams close to $\gamma_{1}=0$. This is shown schematically in Fig. 1 for $\gamma_{2}<0$ and all $|\gamma_{\ell}|$ kept small (see Appendix B). Let us summarize its main aspects [10, 13]:
@@
    At low temperatures and $\gamma_{2}<0$, we obtain an ordered phase with full symmetry breaking (abbreviated f-SSB) and $2Q$ vacua. For $\gamma_{2}>0$, this becomes a first-order line separating partial SSB regions described below (see Appendix B).
@@
The phase diagrams in Fig. 1 are determined by replacing Eq. 1 by an effective gaussian continuum theory [10, 13] via a duality transformation á la Villain [18]
$S\approx\int d^{2}x\left[\frac{\left(\nabla\phi\right)^{2}}{8\pi^{2}\beta}-h\cos(\phi)-\sum_{\ell}\gamma_{\ell}\cos\left(\ell Q\theta\right)\right].$ (5)
and keeping track of the relevance (in the renormalization group sense) of the scaling operators $\cos(\phi),~{}\cos(\ell Q\theta)$ (see Appendix B for details). Recall that a scaling operator $\mathscr{O}$ is relevant when its scaling dimension $[\mathscr{O}]$ is smaller than the spatial dimension (two in our case) for such classical statistical mechanical systems [19]. The term $h\cos(\phi)$ is included as a regulator in the Villain procedure. While the exact relationship between the scaling dimensions and the microscopic parameters $T,h,\gamma_{\ell}$ cannot be determined exactly, close to the gaussian fixed point, we have the following.
$[\cos(\phi)][\cos(\ell Q\theta)]=\ell^{2}Q^{2}/4.$ (6)
Eq. 6 tells us that for $Q\geq 4$, there exists a regime $22>[\cos(Q\theta)]>1/$ 2 where $\cos(Q\theta)$ is the only relevant symmetry-allowed operator. Tuning this away by setting $\gamma_{1}=0$ produces a critical state corresponding to the Landau-incompatible transition or unnecessary critical line. Furthermore, we expect $[\cos(\phi)]$ to increase with $\beta$ to obtain the disordered phase driven by $\cos(\phi)$ at high temperature.
Important parts of the phase diagrams in Fig. 1 have already been explored in previous work (see Refs.[10, 13]). Our main focus will be on symmetry properties, Landau-(in)compatible nature of transitions, unnecessary criticality, and the distinction between even and odd $Q$. These aspects have not previously been investigated, to the best of our knowledge.
Residual symmetries, Landau (in)compatibility: Let us understand the nature of the partial symmetry breaking regions (p-SSB±) which are realized for $\gamma_{1}>0$ and $\gamma_{1}<0$ respectively at intermediate temperatures, where the only relevant operator in Eq. 5 is $\cos(Q\theta)$. Both have $Q$ vacua $\vartheta^{\pm}_{1},\ldots,\vartheta^{\pm}_{Q}$ characterized by the vacuum expectation value $\langle\theta_{j}\rangle=\vartheta^{\pm}_{n}$ with
@@
Since both regions have the same number of vacua, only two possibilities exist: (i) the regions are distinct phases that are Landau-incompatible or (ii) they correspond to the same phase (see Appendix A for more details). To clarify which, we need to determine the residual symmetries of each vacuum $\mathscr{I}\left(\vartheta^{\pm}_{n}\right)$ in both regions. Using Eq. 8, we see that the vacua transform into each other under the discrete rotation symmetry ${\mathcal{R}}$ but preserve a specific $\mathbb{Z}_{2}$ subgroup generated by reflection ${\mathcal{T}}$ followed by a certain number of rotations ${\mathcal{R}}$. To distinguish between various $\mathbb{Z}_{2}$ groups, we define the following notation:
@@
Fig. 1 lists the residual symmetries for representative values of $Q$. For even $Q$, these are distinct for p-SSB± and the vacua are invariant under ${\mathcal{T}}$ followed by even (odd) ${\mathcal{R}}$ rotations for $\gamma_{1}>0~{}(\gamma_{1}<0)$. The p-SSB± phases are detected, respectively, by the following order parameters,
@@
This leads to the conclusion that both belong to the same phase and that there should exist a path where the vacua $\vartheta^{+}_{n+\frac{Q+1}{2}}$ and $\vartheta^{-}_{n}$ can be smoothly connected without encountering a phase transition. This is seen most clearly for $Q=1$ when, in fact, p-SSB± are smoothly connected to the trivial phase and each other (see Appendix B). Thus, for odd $Q$, the critical line along $\gamma_{1}=0$ represents ‘unnecessary criticality’ [11, 12] which does not correspond to a genuine transition separating distinct phases but is nevertheless stable and reached by tuning a single relevant parameter. To the best of our knowledge all known instances of unnecessary criticality [20, 21, 12, 22, 11, 23] have been observed in quantum mechanical systems in a topological ‘failed SPT’ context [24, 25, 26] and ours is the first example in classical symmetry-breaking setting.
@@
For completeness, let us consider the remaining phases and transitions in Fig. 1. The full symmetry breaking phase (f-SSB) appears when $\cos(2Q\theta)$ becomes relevant along the $\gamma_{1}=0$ line at low temperatures, this has $2Q$ vacua characterized by $\langle\theta_{j}\rangle=\chi_{1},\ldots\chi_{2Q}$ with
$\chi_{n}=\frac{\left(2n+1\right)\pi}{2Q},~{}\text{for }n=0,1,\ldots,2Q-1$ (14)
deep in the phase and detected by the order parameter $\mathcal{F}=\sin(Q\theta)$. The symmetries in Eq. 2 transform the vacua into each other as follows (with $\chi_{n}\equiv\chi_{n+2Q}$ identified) as
${\mathcal{R}}:\chi_{n}\mapsto\chi_{n+2},~{}{\mathcal{T}}:\chi_{n}\mapsto\chi_{-n-1}.$ (15)
@@
From Eq. 15, we see that the only residual symmetry for all vacua is the trivial subgroup $\mathscr{I}(\chi_{n})=\{1\}$. It is easy to verify that the transition between p-SSB± and f-SSB is Landau-compatible and belongs to the Ising universality class [17, 14]. Finally, at large values of $T$ along the $\gamma_{1}=0$ line, all symmetries are restored when $\cos(\phi)$ becomes relevant in Eq. 5 to produce a disordered paramagnet (PM). For $Q\leq 4$, the transition between the p-SSB± and the PM is direct and also Landau-compatible whose universality class depends on $Q$ [17, 13] as shown in Fig. 1.
Enhanced symmetries, charged defects:
Figure 3: Two types of domain walls obtained by interpolating the $\langle\theta\rangle=0,\pi$ vacua in the the p-SSB+ phase for even $Q$ (a) hard, neutral domain walls present for $|\gamma_{1}|\rightarrow\infty$ which is invariant under the action of the residual ${\mathcal{T}}$ symmetry. Condensing these restores symmetries (b) soft domain walls obtained for small $|\gamma_{1}|$ which traps a charge and transforms under ${\mathcal{T}}$ which melts to produce DQC. This distinction is exposed by the order parameters $\mathcal{E}_{\pm}$.
@@
We now study two prominent aspects of Landau-incompatible deconfined critical transitions. The first is the appearance of enhanced symmetries involving rotations between order-parameters of the phases that the DQC line separates. This is readily seen for our model. Notice that along $\gamma_{1}=0$, the $D_{2Q}$ symmetry is enhanced to $D_{4Q}$ generated by ${\mathcal{T}}$ and an enhanced $2Q$ fold rotation
~$\tilde{{\mathcal{R}}}:\theta_{j}\mapsto\theta_{j}+{\pi}/{Q}.$ (16)
For even $Q$, this transforms the order parameters for the Landau-incompatible pSSB± phases shown in Eq. 11, into each other.
@@
The second aspect is the physical picture for the onset of DQC being the proliferation of charged defects that prevents the restoration of symmetries [4]. This is also clearly seen in our model. In the vicinity of the DQC transition along $\gamma_{1}=0$, the relevant excitations are smooth interpolations between vacua with the same residual symmetries. The domain walls resulting from this interpolation are charged under the residual $\mathbb{Z}_{2}$ symmetry. For illustration, let us focus on $\gamma>0$ where, for even $Q$, two of the vacua of the resulting p-SSB+ phase are $\langle\theta_{j}\rangle=0\text{ and }\pi$ with the same residual symmetry, $\mathbb{Z}_{2}^{[0]}=\{1,{\mathcal{T}}\}$. If we create a smooth interpolation between these vacua as shown in Fig. 3(b), we see that the resulting domain wall transforms under ${\mathcal{T}}$ and thus carries charge. Furthermore, by evaluating the order parameter $\mathcal{E}_{+}$ on this configuration, we see that it vanishes on the domain wall, whereas the order parameter $\mathcal{E}_{-}$, which vanishes everywhere else, becomes non-zero at the domain wall. Thus, upon melting the p-SSB+ domain walls, we get p-SSB- order!
Other interesting features are revealed by understanding the mechanisms of the other transitions out of the p-SSB phases. In addition to the charged domain walls, we also have ordinary excitations charged under the residual symmetries corresponding to local deviation from the $\gamma_{1}\cos Q\theta$ minima. Condensing these excitations results in further symmetry breaking and the Ising transition into the f-SSB phase. On the other hand, if we go to the $|\gamma_{1}|\rightarrow\infty$ limit, the excitations are ‘hard’ domain walls as shown in Fig. 3(a). These do not transform under the residual symmetries, and their condensation by increasing temperature restores all symmetries. As we reduce $|\gamma_{1}|$, we see that the neutral, hard domain walls soften by absorbing charge, as shown in Fig. 3 and eventually melt to trigger the DQC transition. An analogous charge-binding transition and associated exotic critical phenomena in a 2d classical model were investigated in Refs [27, 28].
@@
Quantum models: All important parts of the phase diagrams of Eq. 1 are qualitatively reproduced by the ground states of the following quantum Hamiltonians
@@
$H_{\text{XXZ}}$ $=\sum_{j}\left(S^{x}_{j}S^{x}_{j+1}+S^{y}_{j}S^{y}_{j+1}+\Delta S^{z}_{j}S^{z}_{j+1}\right),$
$H_{\ell Q}$ $=\sum_{j}\left(\prod_{l=0}^{\ell Q}S^{+}_{j+l}+\prod_{l=0}^{\ell Q}S^{-}_{j+l}\right),$
@@
$\vec{S}=\frac{1}{2}\vec{\sigma}$ are standard spin half angular momentum operators, $H_{\textrm{XXZ}}$ is the XXZ spin chain and $H_{Q}$ is a term that favours SSB. Both choices of $H_{0}$ favour disordered paramagnets that preserve all symmetries, but the second choice also produces a symmetry-protected topological (SPT) phase for one of the signs of $h$. The symmetries in Eq. 2 are generated by
${\mathcal{R}}=\prod_{j}\exp\left(\frac{2\pi i}{Q}S^{z}_{j}\right)$ (18)
@@
whereas ${\mathcal{T}}$ is a time-reversal symmetry generated by complex conjugation in the $Z$ basis. This model can be bosonized to get the same field theory as in Eq. 5 for $|\Delta|<1$, up to renormalization of coupling constants.
The role of anomalies: Deconfined criticality is sometimes associated to the microscopic symmetries being realized anomalously [4] either on the ultraviolet [3, 29, 30, 31] or infrared [8, 9, 32] degrees of freedom  [^1] Note that this prejudice is not universally held, as evidenced by the search for a direct DQC transition between the Néel and valence-bond solid (VBS) phases on the honeycomb lattice [37, 9] which is not constrained by the LSM theorem. However, recent work [40] has indicated that this transition may in fact not exist. Our work proves that even if this particular DQC is not present in the honeycomb lattice, there is no fundamental obstruction to observing Landau-incompatible transitions on the honeycomb or other LSM-trivial lattices.. Anomalies are non-trivial manifestations of symmetries that forbid a strictly on-site representation [34], present an obstruction to gauging [35, 36] and disallow a trivial symmetry-preserving phase [9, 31]. Common settings for anomalies and DQC are systems constrained by Lieb-Schultz-Mattis (LSM) conditions [3, 29, 31, 30] or on the boundaries of symmetry-protected topological phases [34, 7].

### v2

@@ Abstract
Continuous phase transitions where symmetry is spontaneously broken are ubiquitous in physics and often found between ‘Landau-compatible’ phases where residual symmetries of one phase are a subset of the other. However, continuous ‘deconfined quantum critical’ transitions between Landau-incompatible symmetry-breaking phases are known to exist in certain quantum systems, often with anomalous microscopic symmetries. In this Letter, we investigate the need for such special conditions. We show that Landau-incompatible transitions can be found in a family of well-known classical statistical mechanical models with anomaly-free symmetries, introduced by José, Kadanoff, Kirkpatrick and Nelson (Phys. Rev. B 16, 1217). The models are anisotropic deformations of the classical 2d XY model labelled by a positive integer $Q$. For a range of temperatures, even $Q$ models exhibit two Landau-incompatible partial symmetry-breaking phases and a direct transition between them for $Q\geq 4$. Characteristic features of deconfined quantum criticality, such as enhanced symmetries and melting of charged defects are easily seen in a classical setting. For odd $Q$, and corresponding temperature ranges, two regions of a single partial symmetry-breaking phase appear, split by a stable ‘unnecessary critical’ line. We discuss experimental systems that realize these transitions and present anomaly-free quantum models that also exhibit similar phase diagrams.
@@
Spontaneous symmetry breaking (SSB) underpins several important physical phenomena, from the development of long-range orders in matter to endowing mass to fundamental particles [1]. The simplest setting for SSB is when a phase of matter, classical or quantum, with a vacuum invariant under a symmetry group $G$ undergoes a phase transition to produce multiple vacua, each of which preserves only a subset of the original symmetries $H\subset G$. If such a phase transition is continuous, it can be described within the Landau-Ginzburg-Wilson-Fisher (LGWF) framework using a local order parameter field. About twenty years ago, the nature of exotic direct transitions between incompatible SSB quantum phases, where the vacuum symmetries of neither phase could be identified as a subset of the other, was investigated in two-dimensional quantum systems [2]. Although such transitions had appeared in earlier studies [3], Ref. [2] recognized that they could not be framed within the LGWF paradigm in terms of order parameter fields. Instead, they were naturally formulated using gauge fields, which are hidden from sight in the ordered phases but appear at the transition. This prompted the moniker ‘deconfined quantum criticality’ (DQC) [2, 4].
@@
Figure 1: Phase diagrams for the JKKN Hamiltonian shown in Eq. 1 for $Q\geq 2$ and fixed $\gamma_{2}<0$. The arrows represent the different vacua characterized by the expectation value $\langle\theta_{j}\rangle$. The residual symmetry for each vacuum is shown following the notation in Eq. 8. For even $Q\geq 4$, the red line along $\gamma_{1}=0$, represents a direct transition between Landau-incompatible partial symmetry-breaking (p-SSB) phases. For odd $Q\geq 3$, it represents an unnecessary critical transition. The labels Ising, Ashkin-Teller (AT), 3-state Potts, and compact boson ($c=1$) indicate the conformal field theory describing the transition. At high temperatures we have a paramagnetic (PM) phase.
What physical settings can give rise to such Landau-incompatible transitions? Low-dimensional examples [5, 6, 7, 8] and descriptions using non-linear sigma models [9] have clarified that deconfined gauge fields are not essential. However, quantum effects are believed to play a necessary role in DQC [10], especially when viewed from the role played by Berry phases, and most examples of DQC occur under Lieb-Schultz-Mattis conditions [3], when microscopic symmetries are anomalous [11, 12].
@@
In this Letter, we show that these conditions are not necessary and Landau-incompatible transitions can, in fact, be found even in ordinary classical statistical mechanical systems with anomaly-free symmetries. We demonstrate this using a well-known family of models introduced by José, Kadanoff, Kirkpatrick and Nelson (JKKN) [13] obtained by perturbing the 2d classical XY model by on-site anisotropies labelled by a positive integer $Q$. For even $Q\geq 4$, the phase diagram includes a direct phase transition between two Landau-incompatible partial symmetry-breaking phases. This transition displays all notable characteristics of DQC, including the appearance of an enhanced symmetry that rotates between the order parameters of the Landau-incompatible phases, and the melting of charged defects. Interestingly, neutral defects and isolated charges also independently exist and can condense to produce Landau-compatible phase transitions. For odd $Q$, we find an exotic second-order ‘unnecessary critical’ line separating two regions of single phase that does not represent a genuine phase transition [14, 15] but is nevertheless stable. To our knowledge, this is the first case of unnecessary criticality identified in a classical model. We discuss experimental systems that exhibit these phenomena and also present a family of quantum models with anomaly-free symmetries that exhibit similar phase diagrams. Our results show that exotic transitions beyond the LGWF paradigm are more abundant than previously believed.
@@
Models and phase diagram: Let us consider the JKKN model [13], a classical statistical mechanical system of planar rotors $\theta_{j}\sim\theta_{j}+2\pi$ located on the vertices of any two-dimensional lattice. The Hamiltonian is
@@
Here, $j$ labels the vertices and $\langle j,k\rangle$ the edges of the lattice. The relevant symmetry group of the system is generated by a rotation ${\mathcal{R}}$ and a reflection ${\mathcal{T}}$ acting as
@@
and form the non-abelian dihedral group $D_{2Q}\cong{\mathbb{Z}}_{Q}\rtimes\mathbb{Z}_{2}$. We will be interested in the equilibrium phase diagram of Eq. 1, varying the temperature $T=\beta^{-1}$ and $\gamma_{1}$ close to $\gamma_{1}=0$, while keeping all other couplings fixed, the most important being $\gamma_{2}$. This is shown schematically in Fig. 1 for $\gamma_{2}<0$ and all $|\gamma_{\ell}|$ kept small [16]. Let us summarize its main aspects [13, 17]:
@@
    At low temperatures and $\gamma_{2}<0$, we obtain an ordered phase with full symmetry breaking and $2Q$ vacua (abbreviated f-SSB). For $\gamma_{2}>0$, this becomes a first-order line separating partial SSB regions described below [16].
@@
The phase diagrams in Fig. 1 are determined by replacing Eq. 1 by an effective gaussian continuum theory [13, 17] via a duality transformation à la Villain [21]
$S\approx\int d^{2}x\left[\frac{\left(\nabla\phi\right)^{2}}{8\pi^{2}\beta}-h\cos(\phi)-\sum_{\ell}\gamma_{\ell}\cos\left(\ell Q\theta\right)\right],$ (4)
and keeping track of the relevance (in the renormalization group sense) of the scaling operators $\cos(\phi)$ and $\cos(\ell Q\theta)$ [16, 13]. Recall that a scaling operator $\mathscr{O}$ is relevant when its scaling dimension $[\mathscr{O}]$ is smaller than the spatial dimension (two in our case) for such classical statistical mechanical systems [22]. The term $h\cos(\phi)$ is included as a regulator in the Villain procedure. Close to the fixed point described by the conformal field theory (CFT), the scaling dimensions are determined by a single stiffness parameter $K_{\rm eff}$, as [23, 13, 16]
$[\cos(\phi)]=\pi K_{\rm eff},~{}[\cos(\ell Q\theta)]=\frac{\ell^{2}Q^{2}}{4\pi K_{\rm eff}}.$ (5)
While the exact relationship between $K_{\rm eff}$ and the microscopic parameters $T,h,\gamma_{\ell}$ cannot be determined exactly, we see that for $Q\geq 4$, there exists a regime $Q^{2}/(8\pi)<K_{\rm eff}<Q^{2}/(2\pi)$ where $\cos(Q\theta)$ is the only relevant symmetry-allowed operator. Tuning this away by setting $\gamma_{1}=0$ produces a critical state corresponding to the Landau-incompatible transition or unnecessary critical line. Important parts of the phase diagrams in Fig. 1 have already been explored in previous work (see Refs.[13, 17]). Our main focus will be on symmetry properties, Landau-(in)compatible nature of transitions, unnecessary criticality, and the distinction between even and odd $Q$. These aspects have not been investigated previously, to the best of our knowledge.
Residual symmetries, Landau (in)compatibility: Let us understand the nature of the partial symmetry breaking regions (p-SSB±) which are realized for $\gamma_{1}>0$ and $\gamma_{1}<0$, respectively, at intermediate temperatures, where the only relevant operator in Eq. 4 is $\cos(Q\theta)$. Both have $Q$ vacua $\vartheta^{\pm}_{1},\ldots,\vartheta^{\pm}_{Q}$ characterized by the vacuum expectation value $\langle\theta_{j}\rangle=\vartheta^{\pm}_{n}$ with
@@
Since both regions have the same number of vacua, only two possibilities exist: (i) the regions are distinct phases that are Landau-incompatible or (ii) they correspond to the same phase. To clarify which, we need to determine the residual symmetries of each vacuum, $\mathscr{I}\left(\vartheta^{\pm}_{n}\right)$, in both regions. Using Eq. 7, we see that the vacua transform into each other under the discrete rotations ${\mathcal{R}}$ but preserve a specific $\mathbb{Z}_{2}$ subgroup generated by reflection ${\mathcal{T}}$ followed by a certain number of rotations ${\mathcal{R}}$. To distinguish between various $\mathbb{Z}_{2}$ groups, we define the following notation:
@@
Fig. 1 shows the residual symmetries for $Q=2,3,4$. For even $Q$, these are distinct for p-SSB± and the vacua are invariant under ${\mathcal{T}}$ followed by even (odd) ${\mathcal{R}}$ rotations for $\gamma_{1}>0~{}(\gamma_{1}<0)$. The p-SSB± phases are detected, respectively, by the following order parameters,
@@
We conclude that both belong to the same phase and that there should exist a path where the vacua $\vartheta^{+}_{n+\frac{Q+1}{2}}$ and $\vartheta^{-}_{n}$ can be smoothly connected without encountering a phase transition. Examples of explicit paths that connect the two regions are sketched in the end matter. Thus, for odd $Q$, the critical line along $\gamma_{1}=0$ represents ‘unnecessary criticality’ [14, 15] which does not correspond to a genuine transition separating distinct phases but is nevertheless stable and reached by tuning a single relevant parameter. To the best of our knowledge all known instances of unnecessary criticality [24, 25, 15, 26, 14, 27] have been observed in quantum mechanical systems [28, 29, 30], and ours is the first example in a classical SSB setting.
@@
For completeness, let us consider the remaining phases and transitions in Fig. 1. The full symmetry breaking phase (f-SSB) appears when $\cos(2Q\theta)$ becomes relevant along the $\gamma_{1}=0$ line at low temperatures. This is detected by the order parameter
$\mathcal{F}=\sin(Q\theta)$ (13)
@@
and has $2Q$ vacua that break all symmetries. The transition between p-SSB± and f-SSB is Landau-compatible and, when continuous, belongs to the Ising universality class [16, 18]. Finally, at large $T$ all symmetries are restored when $\cos(\phi)$ becomes relevant along the $\gamma_{1}=0$ line to produce a disordered paramagnet (PM). For $Q\leq 4$, this transition is direct and also Landau-compatible; the universality class depends on $Q$ [16, 17].
Figure 2: Schematic representation of various excitations in the pSSB+ phase and their symmetry transformation: (a) local charges, (b) neutral and (c) charged domain walls. The latter can be regarded as a bound state of the former two. The transition triggered by condensing each excitation is shown in (d). The neutral domain wall does not carry symmetry charges and its proliferation restores all symmetries, whereas condensing the charge breaks all symmetries. Melting the charged domain walls produces the exotic transitions discussed in this work: Landau-incompatible transitions for even $Q$ which breaks some symmetries while restoring others, and unnecessary criticality for odd $Q$.
@@
Enhanced symmetries, charged defects: We now study two prominent aspects of DQC transitions in our classical models. The first is the appearance of enhanced symmetries [4] involving rotations between order-parameters of the two Landau-incompatible phases that the DQC line separates. This is readily seen for our model. For even $Q$, we can use the order parameters of p-SSB± phases shown in Eq. 10 to define the two-component unit vector,
$\hat{n}=\begin{pmatrix}\mathcal{E}_{+},\mathcal{E}_{-}\end{pmatrix}=\begin{pmatrix}\cos\left(Q\theta/2\right),\sin\left(Q\theta/2\right)\end{pmatrix}.$ (14)
Along the direct Landau-incompatible transition, the $D_{2Q}$ symmetry of the JKKN model is enhanced to the $O(2)$ symmetry of the XY model whose order parameter is $\hat{n}$. This is generated by a full $\theta$ rotation which transforms the two components of $\hat{n}$, $\mathcal{E}_{\pm}$ into each other.
@@
The second aspect is the physical picture for the onset of DQC being the proliferation of charged defects that prevents the restoration of symmetries [4]. This is also clearly seen in our model. In the vicinity of the DQC transition along $\gamma_{1}=0$, the relevant excitations are smooth interpolations between vacua with the same residual symmetries. The domain walls resulting from this interpolation are charged under the residual $\mathbb{Z}_{2}$ symmetry. For illustration, let us focus on $\gamma_{1}>0$ where, for even $Q$, two of the vacua of the resulting p-SSB+ phase are $\langle\theta_{j}\rangle=0\text{ and }\pi$ with the same residual symmetry, $\mathbb{Z}_{2}^{[0]}=\{1,{\mathcal{T}}\}$. If we create a smooth interpolation between these vacua as shown in Fig. 2(c), we see that the resulting domain wall transforms under ${\mathcal{T}}$ and thus carries charge. Furthermore, by evaluating the order parameter $\mathcal{E}_{+}$ on this configuration, we see that it vanishes on the domain wall, whereas the order parameter $\mathcal{E}_{-}$, which vanishes everywhere else, becomes non-zero at the domain wall. Thus, upon melting the p-SSB+ domain walls, we get p-SSB- order!
Experimental realizations: An outstanding challenge for DQC is the relative paucity of experimental platforms for its validation [4]. The results presented in this Letter open up avenues in classical systems, where Landau-incompatible transitions can be studied more easily. In fact, they have already been observed in several existing experimental systems [31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. For example, the $Q=4$ model of Eq. 1 describes the adsorption of hydrogen on the (100) surface of tungsten [32, 33]. This system exhibits a structural transition [34, 35] which is nothing but the Landau-incompatible transition shown in Fig. 1(c). The same model also describes ultrathin deposits of iron on gold substrate [36, 37, 38, 39].
@@
Quantum models: All important parts of the phase diagrams of Eq. 1 are qualitatively reproduced by the ground states of the quantum Hamiltonian
@@
where, $H_{\text{XXZ}}=\sum_{j}\left(S^{x}_{j}S^{x}_{j+1}+S^{y}_{j}S^{y}_{j+1}+\Delta S^{z}_{j}S^{z}_{j+1}\right),$
$H_{\ell Q}$ $=\sum_{j}\left(\prod_{l=0}^{\ell Q-1}S^{+}_{j+l}+\prod_{l=0}^{\ell Q-1}S^{-}_{j+l}\right),\text{ and }$
@@
$\vec{S}=\frac{1}{2}\vec{\sigma}$ are standard spin half angular momentum operators, $H_{\textrm{XXZ}}$ is the XXZ spin chain and $H_{Q}$ is a term that favours SSB. Both choices of $H_{0}$ favour disordered paramagnets that preserve all symmetries, but the second choice also produces a symmetry-protected topological (SPT) phase for one of the signs of $h$ [20]. The symmetries in Eq. 2 are generated by ${\mathcal{R}}=\prod_{j}\exp({\frac{2\pi i}{Q}S^{z}_{j}})$, ${\mathcal{T}}$ is a time-reversal symmetry generated by complex conjugation in the $Z$ basis and the local order parameters corresponding to Eqs. 10, 11 and 13 are
±$\mathcal{E}_{\pm}=\mathcal{P}_{\pm}(Q/2),~{}\mathcal{O}=\mathcal{P}_{+}((Q-1)/2),~{}\mathcal{F}=\mathcal{P}_{-}(Q)$ (16)
[…]
@@
This model can be bosonized to get the same field theory as in Eq. 4 for $|\Delta|<1$, up to renormalization of coupling constants.
The role of anomalies: The p-SSB± phases have two more distinct type of excitations. The first are charges shown in Fig. 2(a) corresponding to local deviations from the $\gamma_{1}\cos(Q\theta)$ minima, whose condensation further breaks symmetries and produces the transition to the f-SSB phase. The second are ‘hard’ domain walls shown in Fig. 2(b) that do not transform under the residual symmetries and are favoured at large values of $|\gamma_{1}|$. Proliferating these by increasing temperature restores all symmetries and drives the transition to the disordered PM. For small $|\gamma_{1}|$, charges are bound to neutral domain walls to produce the soft domain walls shown in Fig. 2(c). As $\gamma_{1}\rightarrow 0$, these bound states, rather than their constituents, melt to drive the Landau-incompatible transitions studied in this letter.
In several models exhibiting DQC [2, 7, 5], the binding of charges to defects occurs kinematically, due to the anomalous nature of underlying microscopic symmetries [4, 5, 6, 7]. Anomalies are exotic symmetry representations found in systems constrained by Lieb-Schultz-Mattis conditions [3, 41, 42, 43] or on the boundaries of SPT phases [44, 7]. They forbid strictly on-site representations [44], present an obstruction to gauging [45, 46] and disallow a trivial symmetry-preserving phase [12, 43]. The latter feature is reflected in the absence of neutral defects that can proliferate to form a trivial phase and permitting only charged ones which can condense to produce DQC [4].